OpenStudy (anonymous):

When a 10 mL graduated cylinder is filled to the 10 mL mark, the mass of the water was measured to be 9.925 g. If the density of water is taken to be 0.9975 g/mL, what is the percent error for the 10 mL of water

OpenStudy (accessdenied):

Are you having trouble with a particular part of this problem? We need to find the density of water from our mass and volume measured, and then compare it to the actual assumed density given 0.9975 g/mL.

OpenStudy (anonymous):

i can't figure out the whole q

OpenStudy (accessdenied):

The percent error is given as: \( \displaystyle \rm \% \ error = \frac{|Actual - Measured|}{Actual} \times 100\% \) We know the actual density because it is given: 0.9975 g/mL. The density in terms of mass and volume is: Density = Mass / Volume. We have a mass: 9.925 g We have a volume: 10 mL So the density is: Density = 9.925 g / 10 mL which we can simplify. Then it all goes into the % error formula and out pops a percent error: \( \displaystyle \rm \% \ error = \frac{|0.9975 - 0.9925|}{0.9975} \times 100\% \)

OpenStudy (anonymous):

thank you so much

OpenStudy (accessdenied):

No problem! :)